OpenStudy (anonymous):

write the domain and range of using interval notation f(x)=√x-2

OpenStudy (anonymous):

x-2 > =0 x > = 2 x E [2, infinity)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

\[f \left( x \right)=\sqrt{x}-2\] for domain \[ for domain {x}\ge 0 ,D=[0,\infty )\] for range minimum value of f(x)=0-2=-2 \[R=[-2,\infty )\]